# A new long gamma-ray burst formation pathway at solar metallicity

### v1

@@ Acknowledgements.
The POSYDON project is supported primarily by two sources: the Swiss National Science Foundation (PI Fragos, project numbers PP00P2_211006 and CRSII5_213497) and the Gordon and Betty Moore Foundation (PI Kalogera, grant award GBMF8477). MMB was supported by the Boninchi Foundation, the project number CRSII5_21349, and the Swiss Government Excellence Scholarship. OSS acknowledges funding by the INAF Finanziamento per la Ricerca Fondamentale grant no. 1.05.23.04.04. GG and OSS acknowledge financial support by the INAF Grant “GRB PrOmpt Emis- sion Modular Simulator” (POEMS – 1.05.23.06.04) and European Union-Next Generation EU, PRIN 2022 RFF M4C21.1 (PEACE – 202298J7KT). JJA acknowledges support for Program number (JWST-AR-04369.001-A) provided through a grant from the STScI under NASA contract NAS5-03127. PMS, SG, KAR, and MS were supported by the project numbers GBMF8477 and GBMF12341. K.A.R. is also supported by the Riedel Family Fellowship and thanks the LSSTC Data Science Fellowship Program, which is funded by LSSTC, NSF Cybertraining Grant No. 1829740, the Brinson Foundation, and the Moore Foundation; their participation in the program has benefited this work. KK is supported by a fellowship program at the Institute of Space Sciences (ICE-CSIC) funded by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. Z.X. acknowledges support from the China Scholarship Council (CSC). EZ acknowledges support from the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “3rd Call for H.F.R.I. Research Projects to support Post-Doctoral Researchers” (Project No: 7933).
Appendix A Star formation histories
@@
Since the star formation rate density and metallicity evolution impact the cosmic rate density of stable reverse-mass-transfer collapsar LGRB significantly, we perform the same potential energy budget calculations as discussed in Section 2 on the Madau & Fragos (2017) and Neijssel et al. (2019) star formation rate prescriptions. Their cumulative energy distributions are shown in Figures 9 and 10. Because they originate from the same underlying POSYDON stellar populations, the shape of each metallicity distribution is similar, but the star formation rate of each metallicity is redistributed. E.g. $2\leavevmode\nobreak\ \mathrm{Z}_{\odot}$ events have a total cumulative rate of ${\simeq}1000$ Gpc-3 yr-1 in the Madau & Fragos (2017) population, but only ${\simeq}13$ Gpc-3 yr-1 in the Neijssel et al. (2019) population. This is an effect of the latter prescription being fitted to binary black hole mergers, which are predominantly expected to originate from low-metallicity populations. Additionally, their metallicity distribution is narrower compared to Madau & Fragos (2017), which results in less total high-metallicity star formation across the age of the Universe. Figures 12 and 11 show the decrease of the rate density from the energy limits. This effect is similar to that from the Illustris TNG simulation as shown in Figure 5.
@@
Figure 11: The stable reverse mass transfer LGRB rate density from the Madau & Fragos (2017) star formation history.
Figure 12: The stable reverse mass transfer LGRB rate density from the Neijssel et al. (2019) star formation history.
Appendix B Free-fall collapse timescale
@@
These effects would shorten the perceived duration of the collapse compared to the free-fall duration, although the infall via the disk can extend the duration for the fastest rotating models (see, for example Fryer et al. 2024). In general, the free-fall collapse durations of our LGRBs progenitors are of the same order of magnitude as the observed $T_{90}$ values for LGRBs.

### v2

@@ Acknowledgements.
We thank the anonymous reviewer for their insightful comments on the manuscript. The POSYDON project is supported primarily by two sources: the Swiss National Science Foundation (PI Fragos, project numbers PP00P2_211006 and CRSII5_213497) and the Gordon and Betty Moore Foundation (PI Kalogera, grant award GBMF8477). MMB was supported by the Boninchi Foundation, the project number CRSII5_21349, and the Swiss Government Excellence Scholarship. OSS acknowledges funding by the INAF Finanziamento per la Ricerca Fondamentale grant no. 1.05.23.04.04. GG and OSS acknowledge financial support by the INAF Grant “GRB PrOmpt Emis- sion Modular Simulator” (POEMS – 1.05.23.06.04) and European Union-Next Generation EU, PRIN 2022 RFF M4C21.1 (PEACE – 202298J7KT). JJA acknowledges support for Program number (JWST-AR-04369.001-A) provided through a grant from the STScI under NASA contract NAS5-03127. PMS, SG, KAR, and MS were supported by the project numbers GBMF8477 and GBMF12341. K.A.R. is also supported by the Riedel Family Fellowship and thanks the LSSTC Data Science Fellowship Program, which is funded by LSSTC, NSF Cybertraining Grant No. 1829740, the Brinson Foundation, and the Moore Foundation; their participation in the program has benefited this work. KK is supported by a fellowship program at the Institute of Space Sciences (ICE-CSIC) funded by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. Z.X. acknowledges support from the China Scholarship Council (CSC). EZ acknowledges support from the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “3rd Call for H.F.R.I. Research Projects to support Post-Doctoral Researchers” (Project No: 7933).
Appendix A Accretion disk dependence on initial primary mass and period
Within the stable reverse-mass-transfer region, the mass of the accretion disk at collapse depends on the initial primary mass and period. As the initial period or donor mass increase, the accretion disk mass decreases. As such, the lowest mass and shortest period systems within the stable reverse-mass-transfer region, in general, have the most massive accretion disk when collapsing. In this section, we will first cover the initial period dependence and then the initial primary mass dependence. We pick representative slices of the stable reverse-mass-transfer region to show the evolution of different stable reverse-mass-transfer binaries.
A.1 Initial period dependence
At $\mathrm{Z}_{\odot}$, we select systems with $M_{\mathrm{ZAMS,1}}\approx 39.87$ and $q=0.95$, which are undergoing stable reverse mass transfer. These systems are illustrated in Figure 3 and exhibit a correlation between decreasing disk mass and increasing initial period. Only the widest two systems that form an accretion disk at collapse are reversed in disk mass.
In Figure 9, we show the evolution of these models focusing on the mass transfer phases. As expected, the shortest period systems (dark blue) fill their Roche lobe already on the main-sequence, as shown in the third row. They shrink back inside their Roche lobe until core-hydrogen depletion is reached. At this stage, all primaries expand, and for the wider systems (green and yellow), this is their first mass transfer phase. The hydrogen envelopes of the primaries in long-period binaries experience less stripping compared to those in short-period systems; consequently, they exhibit a greater radial expansion after the mass transfer. The $P_{\mathrm{ZAMS}}=37.28\leavevmode\nobreak\ \mathrm{days}$ binary exits the first mass transfer with $4.39\leavevmode\nobreak\ \mathrm{M}_{\odot}$ of hydrogen envelope left and a radius of ${\sim}84R_{\odot}$, while the binary with $P_{\mathrm{ZAMS}}=517.95\leavevmode\nobreak\ \mathrm{days}$ has $8.27\leavevmode\nobreak\ \mathrm{M}_{\odot}$ of hydrogen envelope left and a radius of ${\sim}546R_{\odot}$. This remains until the remaining hydrogen is lost through stellar winds, which also removes all the angular momentum gained during the first mass transfer phase.
Figure 9: The evolution of stable reverse-mass-transfer systems with $M_{\mathrm{ZAMS,1}}\approx 39.87$ and $q=0.95$ and different initial periods, including the first mass transfer and reverse phase. The rows from top to bottom show: the total angular momentum in the primary star, the radius of the primary star, and finally the relative Roche lobe filling factor of the primary and secondary. The colour map indicates the initial period at the zero-age main-sequence, with blue indicating the tightest and yellow the widest systems. The horizontal dashed black lines indicate when the star fills its Roche lobe.
With a mass ratio of $q=0.95$, the secondary evolves off the main-sequence soon after. In the shortest period system, $P_{\mathrm{ZAMS}}\approx 26.83\leavevmode\nobreak\ \mathrm{days}$, the secondary fills its Roche lobe while still on the main sequence, leading to a relatively prolonged and stable mass transfer phase. This results in the transfer of more mass and angular momentum compared to wider-orbit systems, which typically experience their initial reverse interaction upon core hydrogen depletion. In Figure 10, we zoom in on the stable reverse mass transfer and exclude this short-period binary for clarity.
Figure 10: The stable-reverse-mass-transfer phase for the same models as in Figure 9 with the same colour map, except for $P=26.83\leavevmode\nobreak\ \mathrm{days}$ which is not included. The models are aligned when starting the stable reverse-mass-transfer phase. The rows from top to bottom show the total angular momentum in the primary, the relative Roche lobe overflow factor of the secondary, the total hydrogen in the primary, and the fraction of critical rotation at the surface. In the top figure, the shaded areas indicate the reverse mass transfer.
In the top row of Figure 10, we show the total angular momentum in the primary star and the duration of the stable reverse-mass-transfer phase as the shaded region. The shortest period systems (dark blue; ${\sim}12000\leavevmode\nobreak\ \mathrm{yr}$) have a 6 times longer mass transfer than the widest periods (yellow; ${\sim}5000\leavevmode\nobreak\ \mathrm{yr}$). Over time, angular momentum at the surface is able to diffuse inwards to spin up deeper layers in the star. For a longer duration mass transfer, the ongoing mass transfer will replenish the angular momentum of the outer layers of the accretor. As a result, the short-period systems are able to accrete more mass and angular momentum, as the total hydrogen in the third row in Figure 10 shows, despite the surface being near critical rotation for the majority of the mass transfer. The amount of AM per accreted unit mass is constant around $\sim 7$ g cm2 s-1 $\mathrm{M}_{\odot}$. In Figure 11, we show the total amount of angular momentum gained during the stable reverse-mass-transfer phase in blue. The total angular momentum gain in short-period binaries is higher than in long-period systems, but the angular momentum loss in the later evolutionary phases is higher for short-period binaries. Independent of the initial period, approximately 80% of the accreted angular momentum is lost between the stable reverse mass transfer and core carbon depletion.
With the evolutionary state of the primary being similar for different initial periods, the main difference is the duration of the stable reverse mass transfer leading to more mass accretion and angular momentum gain. A long-period binary has a shorter stable reverse-mass-transfer phase, and gains less angular momentum and mass. With a similar fraction of angular momentum lost across different initial periods, this puts most of the long-period primaries outside the regime for disk formation during collapse, while short-period primaries retain sufficient angular momentum to produce an accretion disk, despite the higher absolute angular momentum loss.
Figure 11: The amount of angular momentum (AM) gained during the stable reverse-mass-transfer phase (blue; dots) and the amount of AM lost between the end of the mass transfer and core carbon depletion (red, cross) for the models in Figure 10. The shortest-period systems gain the most angular momentum, but also lose more angular momentum due to their large radial extension compared to long-period binaries. The fractional loss is approximately 80% of the AM gained at each initial period.
A.2 Initial primary mass dependence
The dependence on the initial primary mass is more straightforward compared to the period dependence. At the same period and mass ratio, a more massive primary gains more mass and angular momentum during the stable reverse-mass-transfer phase compared to a lower mass companion, as shown as the blue line in Figure 12 for $P_{\mathrm{ZAMS}}=19.31\leavevmode\nobreak\ \mathrm{days}$ and $q=0.95$ at $Z=\mathrm{Z}_{\odot}$. At the same time, the interaction occurs earlier in the evolution of the primary due to its larger radius. As such, the duration between the reverse mass transfer and core carbon depletion is shorter for lower mass primaries compared to a higher mass primary, despite the longer evolutionary timescale. Together with a smaller radius and lower mass loss through stellar winds, this leads to less angular momentum loss in lower-mass primaries, as the red line in Figure 12 shows. In conclusion, as the initial mass of the primary star increases in a binary, the angular momentum loss before core carbon depletion increases, which decreases the mass of an accretion disk at collapse.
Figure 12: The angular momentum gain (blue) and loss (red) during the reverse mass transfer and post-reverse mass transfer, respectively, as a function of initial donor mass for $P_{\mathrm{ZAMS}}=19.31\leavevmode\nobreak\ \mathrm{days}$ and $q=0.95$ at $Z=\mathrm{Z}_{\odot}$. Only the binaries undergoing stable reverse mass transfer at this period are shown.
Appendix B Star formation histories
@@
Since the star formation rate density and metallicity evolution impact the cosmic rate density of stable reverse-mass-transfer collapsar LGRB significantly, we perform the same potential energy budget calculations as discussed in Section 2 on the Madau & Fragos (2017) and Neijssel et al. (2019) star formation rate prescriptions. Their cumulative energy distributions are shown in Figures 13 and 14. Because they originate from the same underlying POSYDON stellar populations, the shape of each metallicity distribution is similar, but the star formation rate of each metallicity is redistributed. E.g. $\mathrm{Z}_{\odot}$ events have a total cumulative rate of ${\simeq}290$ Gpc-3 yr-1 in the Madau & Fragos (2017) population, but ${\simeq}600$ Gpc-3 yr-1 in the Neijssel et al. (2019) population. This is an effect of the latter prescription having a narrower metallicity distribution compared to Madau & Fragos (2017), which results in more high-metallicity star formation in the local Universe. Figures 16 and 15 show the decrease of the rate density from the energy limits. This effect is similar to that from the Illustris TNG simulation as shown in Figure 5.
@@
Figure 15: The stable reverse-mass-transfer LGRB rate density from the Madau & Fragos (2017) star formation history.
Figure 16: The stable reverse-mass-transfer LGRB rate density from the Neijssel et al. (2019) star formation history.
Appendix C Free-fall collapse timescale
@@
These effects would shorten the perceived duration of the collapse compared to the free-fall duration, although the infall via the disk can extend the duration for the fastest rotating models (see, for example Fryer et al. 2025). In general, the free-fall collapse durations of our LGRBs progenitors are of the same order of magnitude as the observed $T_{90}$ values for LGRBs.